state what you would plot to get a straight line if experimental (x,y) data are to be correlated by the following relations, and what the slopes and intercepts would be in terms of the relation parameters.

for example:

Y^2=ae^(-b/x)

plot: y^2 VS 1/x

Slope: -b

intercept=ln(a)

1- Y^2=mx^3-n

2- 1/ln(y-3)=(1+a.sqrt(x))/b

3- (y+1)^2=[a(x-3)^3]^-1

4- y=e^(a.sqrt(x)+b)

5- xy=10[a(x^2-y^2)+b]

I need

plot:

slope:

intercept:

for all
Answer #1

1.

plot:Y^2 vs X^3

slope:m

intercept:-n

2.

plot:(1/ln(Y-3)) vs sqrt(x)

slope:a/b

intercept:1/b

3.

plot:(Y+1)^2 vs (1/(X-3)^3)

slope:1/a

intercept:0

4.

Y=e^(asqrt(X)+b)

take log both sides

lnY=a*sqrt(X) +b

plot:lnY vs sqrt(x)

slope:a

intercept:b

5.

XY=10((X^2-Y^2)+b)

1/10=a(1/X -1/Y)+b

1/Y=1/X+(1/a)(b-1/10)

plot: 1/Y vs 1/X

slope:1

intercept:(1/a)(b-1/10)
Answer #2

1.

plot:Y^2 vs X^3

slope:m

intercept:-n

2.

plot:(1/ln(Y-3)) vs sqrt(x)

slope:a/b

intercept:1/b

3.

plot:(Y+1)^2 vs (1/(X-3)^3)

slope:1/a

intercept:0